# Multi-objective Representation for Numbers in Clinical Narratives Using CamemBERT-bio

### v2

@@ VI-A Our models
VI-B Baselines
@@
VI-C Setup and Evaluation
The performance evaluation metric utilized is the $F_{1}$ score, a class-wise metric particularly suitable for imbalanced datasets. For the pre-finetuning step, we trained all our models for 100 epochs as a warm up with AdamW optimizer [27] optimizer and a learning rate of $53e-$ 5. Then we used cosine learning scheduler until early-stopping with 4 epochs of patience. For the classification downstream task, we performed 10 sets of runs with different random seeds until early stopping. For all the methods considered in this study, we used cross-entropy loss with appropriate weights for each classes in order to compensate the distribution imbalance. However, such weights were not necessary for our proposed Model 2 which performed better without them. The training, validation, and testing datasets were structured to maintain a distribution of 70%, 15%, and 15% respectively across all eight classes. Our models’ training process took 20 hours on a Tesla V100 GPU with 32 GB of memory. In contrast, training the baseline models required approximately 10 hours. The entire codebase, without the dataset and the trained model weights, can be accessed on our GitHub repository.
VII Results and Discussions
Table  III shows that all models exhibit nearly perfect $F_{1}$ scores for the Out-of-class class. This was expected given its extensive representation. To evaluate overall performance, we compute the average $F_{1}$ score across all classes without applying weighting, considering that the remaining classes hold equal importance in terms of size. This choice is made to ensure a straightforward evaluation without introducing unnecessary complexity. The overall $F_{1}$ scores of all models are presented in Table  I.
TABLE I: Overall $F_{1}$ score per model
@@ VII Results and Discussions
Model 2 0.86
Using this ranking as our basis, it becomes evident that our two models (models 1 and 2) consistently surpass all the other baseline methods.
VII-1 Impact of prefinetuning via MLM
Table II presents a performance comparison of several models with and without prefinetuning. Prefinetuning has no impact on the pretrained CamemBERT-bio model, which is expected since CamemBERT-bio was trained on diverse medical datasets similar to our current dataset, except for the noisy aspect previously explained. However, the performance of the model built on CamemBERT-bio, namely CamemBERT-bio + LESA, significantly improves with prefinetuning. This improvement is attributed to the LESA technique, indicating that incorporating Label Embeddings necessitates updating CamemBERT-bio’s embeddings. Essentially, the latent space derived from CamemBERT-bio + LESA evolves to better discriminate between clusters corresponding to each of the eight classes, thereby enhancing the $F_{1}$ score.
TABLE II: Impact of prefinetuning on $F_{1}$ score
Model Without With
@@ VII-1 Impact of prefinetuning via MLM
VII-2 Impact of Multiple Loss objective
@@
The primary advantage of the number prediction loss function is that it introduces intrinsic magnitude information into the model’s embeddings, an aspect significantly enhanced by incorporating Xval. Through this loss function, we aim for the model to reconstruct numbers with similar magnitudes to the original values. While predicting the exact numbers would lead to memorization, we expect the predictions to be close but not identical. As shown in Figure  5, Model 2 successfully achieves this task.
A detailed examination of the results in Table  III reveals that Model 2 outperforms all other models in nearly every class. Notably, all models struggle with the ventricular Gradient class. This difficulty arises because this class is ambiguous and can be confused with other similar classes, such as aortic gradient, aortic ring, pulmonary gradient, and gradient between earcups. This highlights a limitation of context-based token representation and explains the poor performance of the other models. The range of these gradient values serves as a significant clue for classification, as gradients between earcups are significantly lower than ventricular gradients, for example.
When we subject Model 2 to the same sentence completion test, the results are shown in Figure 6.
@@
This improvement was not observed after prefinetuning CamemBERT-bio, as indicated in Table  II. The performance of CamemBERT-bio remains largely unchanged after prefinetuning. Instead, we attribute the improvement to the implementation of multiple objective number representations. By replacing numbers with placeholders, the text—previously unstructured and filled with abbreviations—becomes clearer, making it easier for the model to learn where numbers are expected. Furthermore, incorporating LESA and Xval enhances the model’s ability to more accurately predict the value range of the masked numbers. We aimed to illustrate our claim by comparing the embeddings of a selection of medical terms generated from various models. Since these medical terms are not numbers, we expect to observe the side effects of multiple objective number representations on the rest of the medical vocabulary. We chose cosine similarity as our metric because, as explained in [28], cosine similarity from token embeddings can be random and meaningless unless the model is trained with respect to cosine similarity, which layer normalization may facilitate. Given that this prerequisite is satisfied in our model, we opted for cosine similarity.
We compared the embeddings generated by the pretrained CamemBERT-bio, the finetuned CamemBERT-bio, Model 1, and Model 2. For consistency, we selected the embeddings from the pretrained CamemBERT-bio as the reference. The cosine similarity graphics are displayed in Figure  7. In this figure, each plot represents a cosine similarity matrix $M$ in grayscale. Each row $i$ and column $j$ is indexed by medical terms $\textrm{tok}_{i}$ and $\textrm{tok}_{j}$, respectively. The corresponding cell $M_{i,j}$ contains the cosine similarity score of the embeddings of $\textrm{tok}_{i}$ obtained by the reference model and the embedding of $\textrm{tok}_{j}$ obtained by the model we’re comparing. The darker a cell is, the higher the cosine similarity score.
The second and third matrices in Figure  7 show that the embeddings generated by the finetuned CamemBERT-bio are significantly different from the reference, while those generated by Model 1 (CamemBERT-bio + LESA) are very similar to the reference. This supports the claim from [2] that simple finetuning significantly disrupts the encoded knowledge of pretrained CamemBERT-bio and strongly overfits, while LESA stabilizes the encoded knowledge of pretrained CamemBERT-bio by guiding the self-attention towards tokens with higher affinity. Regarding the impact of multiple objective number representations, the last matrix in Figure  7 shows that the largest embeddings gap with the reference is observed with Model 2. This evidence supports the claim that integrating multiple losses significantly enhances embeddings compared to using only LESA. Unlike the case of the finetuned CamemBERT-bio in the second matrix, the substantial change in embeddings relative to the reference does not indicate overfitting. This assertion is further validated by the higher F1 score obtained on the test dataset.
Figure 7: Evaluation of the impact of multiple objective numbers representation on the words semantic. From left to right we computed the cosine similarity of multiple medical words embeddings generated by pretrained CamemBERT-bio and itself, pretrained CamemBERT-bio and finetuned CamemBERT-bio, pretrained CamemBERT-bio and Model 1, pretrained CamemBERT-bio and Model 2. The darker a cell is, the higher its value
TABLE III: $F_{1}$ score results per class.
@@
Model 2 0.93 $\pm$ 0.02 0.98 $\pm$ 0.03 0.50 $\pm$ 0.12 0.94 $\pm$ 0.06
VIII Conclusion
@@
This research focused on optimizing numerical understanding by CamemBERT-bio, particularly for classifying numerical values in small medical datasets. We addressed the challenges posed by tokenization and the lack of text structure, which required special handling of numbers. To tackle these issues, we implemented multiple approaches. First, we demonstrated that CamemBERT-bio + LESA could be further improved through prefinetuning with the MLM task, an improvement not seen with CamemBERT-bio alone. Additionally, we developed a multiple-objective numbers representation by combining LESA [2] and Xval [3], resulting in embeddings that incorporate both contextual and magnitude information. Rigorous evaluations of these strategies showed significant enhancements in the performance of CamemBERT-bio.
Acknowledgment
The medical notes utilized in this study were provided by CHU Sainte Justine Hospital and annotated by Dr. Jérôme Rambaud and Dr. Guillaume Sans. The authors express their gratitude to physiotherapist Kevin Albert for his contribution to the literature review on medical benchmarks.
This research was partially supported by the Natural Sciences and Engineering Research Council (NSERC), the Institut de Valorisation des Données de l’Université de Montréal (IVADO), and the Fonds de la Recherche en Santé du Québec (FRQS).

### v3

@@ 6.1 Our models
6.2 Baselines
GPT-4: In these experiments, we selected a subset of 200 entries from our dataset and manually provided them to ChatGPT using a few-shot learning prompting approach. The model generated responses as sequences of numerical values, each mapped to its respective category. These outputs were then aggregated and used to compute the evaluation metrics detailed below.
@@ 6.2 Baselines
6.3 Setup and Evaluation
The performance evaluation metric utilized is the $F_{1}$ score, a class-wise metric particularly suitable for imbalanced datasets. For the pre-finetuning step, we trained all our models for 100 epochs as a warm up with AdamW optimizer [31] optimizer and a learning rate of $53e-$ 5. Then we used cosine learning scheduler until early-stopping with 4 epochs of patience. For the classification downstream task, we performed 10 sets of runs with different random seeds until early stopping. We compute the mean F1 score over the 10 runs along with their standard deviations, with all results summarized in Table  3. For all the methods considered in this study, we used cross-entropy loss with appropriate weights for each classes in order to compensate the distribution imbalance. However, such weights were not necessary for our proposed Model 2 which performed better without them. The training, validation, and testing datasets were structured to maintain a distribution of 70%, 15%, and 15% respectively across all eight classes. Our models’ training process took 20 hours on a Tesla V100 GPU with 32 GB of memory. In contrast, training the baseline models required approximately 10 hours.
7 Results and Discussions
Table  3 shows that all models exhibit nearly perfect $F_{1}$ scores for the Out-of-class class. This was expected given its extensive representation. To evaluate overall performance, we compute the average $F_{1}$ score across all classes without applying weighting, considering that the remaining classes hold equal importance in terms of size. This choice is made to ensure a straightforward evaluation without introducing unnecessary complexity. The overall $F_{1}$ scores of all models are presented in Table  1.
Table 1: Overall $F_{1}$ score per model
@@ 7 Results and Discussions
Model 2 0.90
Using this ranking as our basis, it becomes evident that our two models (models 1 and 2) consistently surpass all the other baseline methods except for GPT-4
7.0.1 IMPACT of PRE-FINETUNING via MLM
Table 2 presents a performance comparison of CamemBERT-bio and CamemBERT-bio + LESA before and after prefinetuning. The results show that Prefinetuning has no significant impact on CamemBERT-bio model, which is expected since CamemBERT-bio was already trained on diverse medical datasets similar to our current dataset, which means that most of the medical knowledge was already encoded in its parameters. However, the results show a significant performance boost for CamemBERT-bio + LESA after prefinetuning. This improvement is attributed to the LESA technique and it indicates that incorporating Label Embeddings in the model’s architecture requires fine-tuning in order to adjust the parameters of CamemBERT-bio’s embedding layers. Essentially, the latent space derived from CamemBERT-bio + LESA evolves to better discriminate the eight categories, thereby enhancing the $F_{1}$ score.
Table 2: Impact of prefinetuning on $F_{1}$ score
Model Before After
@@ 7.0.1 IMPACT of PRE-FINETUNING via MLM
7.0.2 RESULTS of MULTIPLE OBJECTIVE LOSS
@@
The primary advantage of the number prediction loss function is that it introduces intrinsic magnitude information into the model’s embeddings, an aspect significantly enhanced by incorporating Xval. Through this loss function, we aim for the model to reconstruct numbers with similar magnitudes to the original values. While predicting the exact numbers would lead to memorization, we expect the predictions to be close but not identical. As shown in Figure  5, Model 2 the model predictions on the $y$ axis mostly aligns with the ground truth numbers on the $x$ axis. This shows that the goal was successfully achieved. We did not include the mean square error, as it is not a key metric in our evaluation. Since our goal is not to have predictions overly close to the ground truth, emphasizing this metric could encourage the model to memorize rather than generalize.
Table  3 presents the F1 score performance of all studied models, broken down by category. For each category, the highest F1 score achieved by a model, excluding GPT-4, is highlighted in bold. A detailed examination of the results in Table  3 reveals that Model 2 outperforms all other models except GPT-4 in nearly every class. Notably, all models, including GPT-4 struggle with the ventricular Gradient class. This difficulty arises because this class is ambiguous and can be confused with other similar classes, such as aortic gradient, aortic ring, pulmonary gradient, and gradient between the heart earcups. This highlights a limitation of context-based token representation and explains the poor performance of the other models. Indeed, in certain failure cases, the surrounding context provides insufficient cues to accurately determine the gradient type. Instead, the key distinguishing factor lies in the magnitude of the gradient values. For example, gradients between heart earcups are notably lower than ventricular gradients. This magnitude-based distinction gives Model 2 a significant advantage over the other models, resulting in a higher F1 score for the gradient category.
When we subject Model 2 to the same sentence completion test as in section  4.1, the results are shown in Figure 6.
@@
Figure 7: Evaluation of the impact of Multiple Objective Representation on the words embeddings. From left to right we computed the cosine similarity of multiple medical terms embeddings generated by CamemBERT-bio and itself, finetuned CamemBERT-bio, Model 1 and Model 2 respectively. The darker a cell is, the higher its value
Table 3: $F_{1}$ score results per class.
@@
Model 2 0.93 $\pm$ 0.02 0.98 $\pm$ 0.03 0.61 $\pm$ 0.12 0.94 $\pm$ 0.06
The side-by-side comparison of Figure  2 and  6 is highly illustrative. Unlike CamemBERT-bio, Model 2 clearly recognizes that a numerical value is expected in this context. Moreover, when we take the prediction from the Number Head layer, we obtain $13.9$, which falls within the correct range for this type of gradient. We attribute this improvement to the implementation of multiple objective number representations. By replacing numbers with placeholders, the unstructured and noisy text becomes clearer, making it easier for the model to learn where numbers are expected. Furthermore, incorporating LESA and Xval enhances the model’s ability to more accurately predict the value range of the masked numbers. We aim to illustrate our claim by measuring the similarity of the embeddings vectors produced by various models derived from CamemBERT-bio and those generated by CamemBERT-bio itself (our reference). Specifically, we evaluate three derived models: CamemBERT-bio fine-tuned on unannotated medical notes using masked language modeling (MLM), Model 1, and Model 2. This experiment is designed to assess how the different methodologies developed in this study influence the embedding space of CamemBERT-bio. To conduct this analysis, we selected twenty frequently occurring medical terms from our dataset. As a similarity metric, we employed cosine similarity, following the insights from [32], which highlights that cosine similarity is a meaningful measure primarily for models trained with respect to this metric—a condition that can be facilitated by layer normalization. Since our models satisfy this prerequisite, cosine similarity was chosen as the most appropriate metric for our evaluation.
The cosine similarity graphics are displayed in Figure  7. In this figure, each plot represents a cosine similarity matrix $M$ in grayscale. Each row $i$ and column $j$ is indexed by the medical terms $\textrm{tok}_{i}$ and $\textrm{tok}_{j}$, respectively. The corresponding cell $M_{i,j}$ contains the cosine similarity score of the embeddings of $\textrm{tok}_{i}$ obtained by the reference model and the embedding of $\textrm{tok}_{j}$ produced by the model we’re comparing. The darker a cell is, the higher the cosine similarity score.
From left to right, the matrices progressively lighten, indicating increasing divergence of the embedding vectors from the reference model. Specifically, the second matrix demonstrates that CamemBERT-bio embeddings remain largely unchanged after fine-tuning, corroborating the results in Table  2, which suggest that fine-tuning alone does not introduce additional knowledge into the model. In contrast, the third matrix reveals a notable difference in embeddings between Model 1 and CamemBERT-bio. This observation aligns with the analysis in Table  2, which suggests that fine-tuning Model 1 via MLM modifies the embedding space due to the integration of label information into CamemBERT-bio through the LESA technique. Finally, the last matrix in Figure  7 exhibits the most substantial deviation from the reference, highlighting the impact of Multiple Objective Representations. This result supports the hypothesis that incorporating multiple losses leads to a more significant transformation of the embedding space compared to using LESA alone.
7.1 GPT-4 VS MODEL 2
The category-wise and overall F1 score analysis highlights GPT-4’s superior performance, with an overall F1 score surpassing Model 2 by 0.05 units. However, this improvement is relatively minor given the vast disparity in model size—Model 2, with 110 million parameters, compared to GPT-4’s multi-trillion parameter scale.
A failure case analysis conducted by [2] identified recurring misclassifications by GPT-4, particularly in the Contractibility (Cp) and Gradient (G) categories. Misclassifications in the Cp category often stemmed from ambiguous abbreviations, whereas errors in the G category frequently involved confusion between different gradient types, such as aortic and atrial/ventricular gradients, which were mistakenly linked to the interventricular gradient. These systematic errors suggest limitations in GPT-4’s contextual comprehension of medical terminology. Similarly, Table  3 shows that Model 2 exhibits difficulties in the Cp and G categories. However, it significantly reduced the performance gap compared to the difference between CamemBERT-bio and GPT-4, narrowing it from 0.52 to 0.17 F1 score units in the G category.
From a practical perspective, our approach offers notable advantages in terms of computational efficiency and system integration. Unlike large-scale LLMs such as GPT-4 and LLaMA, Model 2 can be seamlessly deployed in medical decision-support systems while maintaining competitive performance. Its reduced computational demands enable local execution, eliminating reliance on external cloud-based models and thereby enhancing data privacy and accessibility.
8 Limitations and Application Perspectives
The main limitation of this study is the choice of CamemBERT-bio over large language models (LLMs). In a context where scalable models are emphasized, this decision was driven by the need to balance performance with computational efficiency. While a direct comparison with GPT-4 confirms its superior overall performance, its computational demands are significantly higher, making it less practical for integration into resource-constrained environments.
Additionally, we recognize that our experiments focus on a specific task and rely on a single dataset. However, we believe that the methodologies introduced in this work are broadly applicable and can be adapted to similar scenarios involving limited data availability.
The Clinical Decision Support System (CDSS) at CHUSJ is still under development. Our next step is to integrate this NLP model for clinical text analysis with previously developed algorithms for heart failure detection [33, 34, 35], hypoxemia detection [36], and chest X-ray analysis [37, 38]. Ultimately, we aim to deploy a comprehensive CDSS within the pediatric intensive care unit (PICU) at CHUSJ to facilitate early diagnosis of acute respiratory distress syndrome (ARDS). Once integrated into the PICU’s e-Medical infrastructure, we will conduct a prospective evaluation of the system’s ARDS detection capabilities, using the results to guide further refinements and enhancements.
9 Conclusion
@@
This research focused on optimizing numerical understanding by CamemBERT-bio, particularly for classifying numerical values in small medical datasets. We addressed the challenges posed by tokenization and the lack of text structure, which required special handling of numbers. To tackle these issues, we implemented multiple approaches. First, we demonstrated that CamemBERT-bio + LESA could be further improved through prefinetuning with the MLM task, an improvement not seen with CamemBERT-bio alone. Additionally, we developed a multiple-objective numbers representation by combining LESA [2] and Xval [3], resulting in embeddings that incorporate both contextual and magnitude information. Rigorous evaluations of these strategies showed significant enhancements in the performance of CamemBERT-bio. While our results do not surpass those of GPT-4, they remain comparable despite the significant difference in model size. This demonstrates the effectiveness of our approach in achieving a balance between performance and computational efficiency. We believe that this work presents a promising alternative to large-scale LLMs for medical note processing, particularly in hospital environments where computational resources are limited and data availability constraints must be considered.